OpenStudy (anonymous):

Evaluate the integral of sin x/2 dx. I know this should be easy, but my answer isn't the same as the answer key. is the answer really -2cosx/2 + C?

OpenStudy (anonymous):

is that sin (x/2) or (sin x) / 2

OpenStudy (anonymous):

the answer you provided is correct if the question is \[\int\limits_{}^{} \sin \frac{ x }{ 2 } dx\]

OpenStudy (anonymous):

oh, could you explain to me where the -2 in front of cos comes from? I was sort of confused by that.

OpenStudy (anonymous):

the negative part, or the 2? the negative comes from the fact that integration of sin = - cos the 2 comes from the substitution u = x/2, du = 1/2 dx, and so dx = 2 du